OpenStudy (anonymous):

) The acceleration due to gravity, g, varies with height above the surface of the earth, in a certain way. If you go down below the surface of the earth, g varies in a different way. It can be shown that g is given by Here R is the radius of the earth, M is the mass of the earth, G is the gravitational constant, and r is the distance to the center of the earth. Sketch a graph of g against r, and use it to answer the following questions: A. Is g a continuous function of r? why? B. Is g a differentiable function of r? why?

OpenStudy (goformit100):

Please close the question. Post the Question in the correct Section.

OpenStudy (goformit100):

Anyways, A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (ybarrap):

When \(R<R_e\), \(g=\cfrac{GM}{R^2}=\cfrac{4\pi R^3\rho}{R^2}=\cfrac{4}{3}G\pi R \rho\), where \(\rho\) is the density of the earth, \(G\) is the gravitational constant, \(R_e\) is the radius of the earth and \(M\) is the mass of the earth. In this case, \(g\) varies \(linearly\) with \(R\). When \(R>R_e\), \(g=\cfrac{GM}{R^2}\) and \(g\) varies in an inverse square fashion with \(R\):  You can check if both are continuous and differentiable at \(R=R_e\), which would be the only points of concern. Use the definition of \(limit\).